### v1

@@ 1 Introduction
The Eigenspace Perturbation Framework (EPF) tries to estimate the model-form uncertainty of the the turbulence model and helps optimize design with reduced sensitivity to uncertainty Emory et al. (2013); Iaccarino et al. (2017); Mishra & Iaccarino (2019). It has been successfully applied in aerospace, turbomachinery, civil structural, and wind farm design Mishra & Iaccarino (2017); Cook et al. (2019); Mishra et al. (2020; 2019); Lamberti & Gorlé (2019); Eidi et al. (2021). Data-driven approaches, aided by high-fidelity simulations and Machine Learning (ML), are gaining popularity in RANS turbulence modeling Heyse et al. (2021); Matha & Kucharczyk (2022); Matha et al. (2023). This study focuses on the use of data-driven methods in order to identify flow regions with potential turbulence model prediction inaccuracies. Random forests (RF) are trained and validated using additional data from scale-resolving simulations. We analyze a methodology to assess a priori prediction confidence. Our framework is an amalgam of physics based domain knowledge with data-driven RF.
@@ 2 Eigenspace Perturbation Framework
as illustrated in Figure 1. The perturbation strength $p$ should be predicted by a ML model in order to modify the Reynolds stress towards the same three limiting states as in the physics-constrained data-free approach. Hereby, on top of the foundation of the EPF, which is allowing for understandable uncertainty estimates for turbulence modeling, we incorporate data-driven ML techniques to enhance the predictive capabilities of the entire framework. The final application of propagating these perturbations throughout the RANS solver is beyond the scope of this paper (for details on that see Heyse et al. (2021); Matha et al. (2023)).
@@ 3.3 Verification of ML model
The prediction accuracy of the RF should be evaluated on the available data (see Section 3.2) with fitted hyperparameters (maximum tree depth = 15, minimum sample count = 10, maximum number of features = 7, number of trees = 30) based on leave-one-out-cross validation (not shown here). Ten different scenarios combining individual flow cases, listed in Table 3, serve as verification of functionality and present the accomplishment of the intended generalization of every model. Since the target quantity $p$ ranges between zero and one (due to the construction of the equilateral triangle with an edge length of one), the resulting $\mathrm{RMSE}$ values indicate less than 10% absolute prediction error, except for scenario II.
@@ 4 Conclusion and Outlook
Our approach represents a hybrid framework that combines a physics-based UQ methodology with data-driven ML. Using ML predicted, spatially varying perturbation strength helps CFD practitioners to identify regions, in which the turbulence model might be erroneous on the one hand. Propagating these prediction to the RANS solver will help to obtain data-driven, less conservative and nonetheless physics-constrained uncertainty estimates for QoI.

### v2

@@ 1 Introduction
The Eigenspace Perturbation Framework (EPF) tries to estimate the model-form uncertainty of the the turbulence model and helps optimize design with reduced sensitivity to uncertainty Emory et al. (2013); Iaccarino et al. (2017); Mishra & Iaccarino (2019). It has been successfully applied in aerospace, turbomachinery, civil structural, and wind farm design Mishra & Iaccarino (2017); Cook et al. (2019); Mishra et al. (2020; 2019); Lamberti & Gorlé (2019); Eidi et al. (2021). Data-driven approaches, aided by high-fidelity simulations and Machine Learning (ML), are gaining popularity in RANS turbulence modeling Heyse et al. (2021); Matha & Kucharczyk (2022). The present study, which was the foundation of the CFD application results in our previous paper Matha et al. (2023), focuses on the use of data-driven methods in order to identify flow regions with potential turbulence model prediction inaccuracies. Random forests (RF) are trained and validated using additional data from scale-resolving simulations. We analyze a methodology to assess a priori prediction confidence. Our framework is an amalgam of physics based domain knowledge with data-driven RF.
@@ 2 Eigenspace Perturbation Framework
as illustrated in Figure 1. The perturbation strength $p$ should be predicted by a ML model in order to modify the Reynolds stress towards the same three limiting states as in the physics-constrained data-free approach. Hereby, on top of the foundation of the EPF, which is allowing for understandable uncertainty estimates for turbulence modeling, we incorporate data-driven ML techniques to enhance the predictive capabilities of the entire framework. The final application of propagating these perturbations throughout the RANS solver is beyond the scope of this paper (for details on that see Heyse et al. (2021); Matha et al. (2023)). Additionally, ML predicted $p$, representing the deviation of barycentric coordinates from high-fidelity data to RANS simulations, might be helpful for future turbulence modeling activities aiming for correct representation of Reynolds stress anisotropy Singh et al. (2017).
@@ 3.3 Verification of ML model
The prediction accuracy of the RF should be evaluated on the available data (see Section 3.2) with fitted hyperparameters (maximum tree depth = 15, minimum sample count = 10, maximum number of features = 7, number of trees = 30) based on leave-one-out-cross validation (not shown here). Ten different scenarios ,combining individual flow cases listed in Table 3 for training and testing purposes, serve as verification of functionality and present the accomplishment of the intended generalization of every model. Since the target quantity $p$ ranges between zero and one (due to the construction of the equilateral triangle with an edge length of one), the resulting $\mathrm{RMSE}$ values indicate less than 10% absolute prediction error, except for scenario II.
@@ 4 Conclusion and Outlook
Our approach represents a hybrid framework that combines a physics-based UQ methodology with data-driven ML. Using ML predicted, spatially varying perturbation strength helps CFD practitioners to identify regions, in which the turbulence model might be erroneous on the one hand. Propagating these prediction to the RANS solver helps to obtain data-driven, less conservative and nonetheless physics-constrained uncertainty estimates for QoI on the other handMatha et al. (2023).